1. A 35 πœ‡πΆ point charge is placed 32 π‘π‘š from an identical 32 πœ‡πΆ charge. How much work would be required to move a 0.50 πœ‡πΆ charge from a point midway between them to a point 12 π‘π‘š closer to either of the charges?


1
Expert's answer
2021-04-29T10:41:55-0400

Explanations & Calculations


  • Work is the energy needed to move between the two mentioned points and it is given by

"\qquad\qquad
\begin{aligned}
\small W&=\small q\Delta V
\end{aligned}"

  • Now q is known & the potential difference between those points are to found to complete the math.
  • Potential of the point midway between the like charges is

"\qquad\qquad
\begin{aligned}
\small V_1&=\small k\frac{q_1}{r_1}+k\frac{q_2}{r_2}\\
&=\small (9\times 10^9Nm^2C^-2 )\Bigg[\frac{35\times 10^{-6}C}{0.16m}+\frac{35\times10^{-6} C}{0.16m}\Bigg]\\
&=\small 3.9375\times 10^6 V
\end{aligned}"

  • Potential of the point 12m from each charge

"\qquad\qquad
\begin{aligned}
\small V_2 &=\small (9\times10^9)\Bigg[\frac{35\times 10^-6}{0.12}+\frac{35\times 10^{-6}}{0.12}\Bigg]\\
&=\small 5.25\times10^6V
\end{aligned}"

  • Then the required work is

"\qquad\qquad
\begin{aligned}
\small W&=\small 0.50\times 10^{-6}C\times 5.25\times 10^6V\\
&=\small \bold{2.625J}
\end{aligned}"


  • The "\small 0.5\mu C" charge moves perpendicular to both charges